OpenStudy (anonymous):

The Law of Cosines states that a^2 = b^2 - c^2 + 2bc cos A. -Is this true or false? I can't find the answer anywhere in my book.

OpenStudy (zarkon):

google law of cosines

OpenStudy (johnweldon1993):

False I believe it should look like a² = b² + c² - 2bc cosA

OpenStudy (anonymous):

the 2bc should be negative but besides that it is correct

OpenStudy (anonymous):

so i guess technically it is false

OpenStudy (anonymous):

it false. Thanks for the help guys!

OpenStudy (johnweldon1993):

no problem!

OpenStudy (anonymous):

\[\cos A=\frac{ b ^{2}+c ^{2}-a ^{2} }{2bc }\] \[2bc \cos A=b ^{2}+c ^{2}-a ^{2}\] now you can solve